The upper Angel Falls, the highest waterfall on Earth, are 750 m higher than Niagara Falls. If each of the falls were 7m lower, the upper Angel Falls would be 16 times as high as Niagara Falls. How high is each waterfall?
Answer by ankor@dixie-net.com(12701)

"The upper Angel Falls, the highest waterfall on Earth, are 750 m higher than Niagara Falls."
Let x = height of Niagara Falls, then (x+750) = height of Angel falls
:
"If each of the falls were 7m lower, the upper Angel Falls would be 16 times as high as Niagara Falls."
16(x-7) = (x+750) - 7
16x - 112 = x + 743
16x - x = 743 + 112
15x = 855
x = 855/15
x = 57
:
How high is each waterfall?
Niagara height = 57 ft
Angels height = 750 + 57 = 807 ft
:
:
Check solution using this statement:
:If each of the falls were 7m lower, the upper Angel Falls would be 16 times as high as Niagara Falls."
807 - 7 = 16(57 -7)
800 = 16(50)